Fermi National Laboratory
 

Accelerator Update for October 23, 2007 - October 24, 2007

 

Tuesday October 23

The day shift began with the accelerator complex in access.

Linac experts established beam to the Linac absorber at 11:34 AM. Linac Work: they fixed the Debuncher, connected an NMR probe to the spectrometer magnet, replaced failed counters on a Klystron RF station (KRF3), and checked the brushes on the I- Source Haefely.

The Tevatron (TeV) came out of access at 1:49 PM.

The Antiproton Source (Pbar) came out of access at 2:41 PM.

Operations reported at 3:09 PM that the SwitchYard 120 (SY) and the Meson primary beamlines have been given permission to take beam.

Operators began a search and secure of the Collider Detector at Fermilab (CDF) at 3:09 PM.

Operations put current on the TeV bus at 4 PM.

Operators completed a search and secure of D-Zero (D0) at 5:39 PM. The D0 experiment successfully completed their preamp repair.

Operations established a Main Injector (MI) ramp at 8 PM.

Operations put the TeV through a dry squeeze at 8:12 PM.

Operations established beam to the MI absorber at 8:29 PM.

TeV experts began a study period around 8:47 PM.

Pbar resumed stacking around 11 PM.

Wednesday October 24

The midnight shift began with Pbar stacking (13.5mA), with the Recycler and TeV experts conducting studies, and with SciBooNE / MiniBooNE taking beam.

The Recycler expert studies ended around midnight. However, operators completed a requested Recycler study during the shift.

At 1:10 AM, the MI power supply interlock loop tripped off. It reset.

Operations prepared the Recycler to stash antiprotons at 2:25 AM.

Around 6:14 AM, the MIPS interlock loop tripped off for the first of two times. Operators reset it.

At 6:38 AM, a Pbar power supply (M:V105) tripped off on a magnet over temperature indication. Operators contacted an expert and per his instructions reset the supply.

Operations reported that the Pbar stack rate had been poor all shift. At 7:21 AM, Operators turned off the Pbar momentum cooling system; the stack rate improved.

At 7:42 AM, the Pbar beam permit dropped on a “Target Station Permit – Dump.” An expert investigated, but wasn’t able to reset the trip. Another expert investigated.

Operations put Pbar into standby mode at 7:53 AM for the investigation.

The Plans for Wednesday

The current plan for today is for the accelerators to continue with startup commissioning.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/24/2007    email Fermilab